OpenStudy (anonymous):

Recall that the equation for horizontal distance "h" in feet of a projectile with initial velocity v0 and initial angle theta is given by h=(v0^2/16)sin theta cos theta. a.) Assume the initial velocity is 60 (feet/second). What initial angle will you need to ensure that the horizontal distance will be exactly 100 feet? b.) given an initial angle of .6 radians, what velocity will be needed to ensure that the horizontal distance will be exactly 100 feet?

OpenStudy (marco26):

You all the required parameters in a and b to solve the question. Just substitute them and find the missing one.

OpenStudy (anonymous):

I have \[\frac{ 60^2 }{ 16 }\sin \theta \cos \theta=100\] and then it will equal sin 2 theta=.8889. This will give you theta equal to 0.547 radians

OpenStudy (anonymous):

That is for part A

OpenStudy (marco26):

Yes. You're right.

OpenStudy (anonymous):

What about part B?